# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

### v2

@@ VI Conclusion
In this paper, we introduce PennyLang, a high-quality, instruction-response dataset tailored for PennyLane-based quantum code generation. We propose a robust data collection and annotation methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples. To assess the practical utility of the dataset, we develop a Retrieval-Augmented Generation (RAG) evaluation framework and benchmark three state-of-the-art LLMs—GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B—under both RAG and non-RAG conditions. Our findings demonstrate consistent performance gains with RAG, especially in functionality and syntax correctness. Building on this, we introduce a GraphRAG extension that models structural relationships across code snippets using graph-based retrieval. When applied to GPT-4o Mini, GraphRAG yields a 2.8$\times$ improvement in accuracy compared to the non-RAG baseline, highlighting its potential for deeper semantic retrieval in quantum programming. We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development. Future directions include training lightweight instruction-tuned models on this dataset, scaling GraphRAG across different models and quantum frameworks, and expanding toward cross-framework generalization between PennyLane, Qiskit, and Cirq.

### v3

@@ VI Conclusion
In this paper, we introduce PennyLang, a high-quality instruction–response dataset tailored for PennyLane-based quantum code generation. We present our data collection and prepossessing methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples. To evaluate its practical utility, we develop a RAG evaluation framework and benchmark four state-of-the-art LLMs (GPT-4o Mini, Claude 3.5 Sonnet, Qwen2.5-7B, and LLaMa 4-17B Maverick) under both RAG and non-RAG settings. Our findings show that PennyLang, when paired with RAG, substantially boosts performance for open-source models, while proprietary models exhibit minimal or negative gains with naive full-context retrieval, likely because their pre-training already covers overlapping PennyLane-related knowledge. This further highlights the importance of calibrated retrieval budgets, as moderate settings (e.g., 75%) often outperform full context by balancing signal and noise. We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development. Future directions include training lightweight instruction-tuned models on this dataset, optimizing retrieval strategies across different model types, and extending to cross-framework generalization between PennyLane, Qiskit, and Cirq.